# MS-XXXX-XXXX.XX

### v1

@@ 6.1 Cuts for WE
where $x^{f}$ is a fixed fractional solution to WDP-LP, treated here as a constant, as is $\bm{D}$. (If WDP-LP admits an integral solution, we are at a WE, and so we would terminate without the need for a new cut.) While constraints (10) ensure that $(\alpha,\alpha_{0})$ is a valid cut for WDP, any positive scalar multiple is also valid, resulting in an unbounded objective. To avoid this unboundedness, one can begin with $\alpha_{0}=1$ and later scale ($\alpha,\alpha_{0}$) to find an integer cut, as described in Algorithm 1.
Algorithm 1 Heuristic to Find AWE Prices
Solve WDP and WDP-LP
     Solve (cut-frac) with $\bar{\alpha}\in\mathbb{R}^{K}_{\geq 0}$ and fixed $\bar{\alpha}_{0}=1$
     Find the smallest $\gamma\geq 1$ such that $\forall k\in\mathcal{K}$, $\gamma\bar{\alpha}_{k}\in\mathbb{Z}$
     Update $(\alpha,\alpha_{0})\leftarrow(\gamma\bar{\alpha},\gamma\bar{\alpha}_{0})$
     Add cut $(\alpha,\alpha_{0})$ to all formulations
Solve WDP-quad-dual (including all cuts as new artificial items)
Proposition 6.1
Algorithm 1 converges to AWE prices for any CA instance.
Solved as an LP (i.e., with no integrality constraints within the while loop), Algorithm 1’s cut generation appears easy; the difficulty is in generating the large matrix $\bm{D}$, requiring the enumeration of solutions to WDP, which is NP-Hard in general. For small examples, this is not problematic, but as the auction becomes larger, columns of $\bm{D}$ will need to be dynamically generated to manage the size. (Details on the dynamic generation of $\bm{D}$ are described in Appendix 8.) Alternatively, since it is clear from the related literature that there are often many ways to find cuts to a non-integral IP formulation, we provide Algorithm 1 as just one direct cutting algorithm for completeness, leaving the development of faster or more specialized techniques for future research.
@@ 6.2 Cuts for any Core Point and PME
Example 4.2 demonstrated that a WE may not be a PME, motivating us to consider additional cuts to WDP to get to PME. Algorithm 1 can take us no further in generating new cuts once we have an AWE (or if we start at NWE), as its progress was driven by closing an IP gap that is now closed. For another, perhaps more interesting illustration, consider the single-minded bidders in the Example 6.2 that follows. Here there are $c_{A}=17$ copies of one item, $A$. (This problem is adapted from an example on lifting to get cuts to the knapsack problem from Bertsimas and Weismantel 2005.)
@@
Proof 6.4
$\min\quad$ $\alpha\bm{0}$
s.t. $pa^{k}+\alpha_{k}=\bar{\rho}_{i(k)}$ $\forall k\in\mathcal{K}^{*}$
$pa^{k}+\alpha_{k}\geq b^{k}-b^{i(k)*}+\bar{\rho}_{i(k)}$ $\forall k\in\mathcal{K}\setminus\mathcal{K}^{*}$
$\alpha\bm{D}\geq\bm{0}$
$p\in\mathbb{R}^{J}_{\geq 0}\text{ ,}\quad\alpha\in\mathbb{R}^{K}_{\geq 0}$
By construction, the existence of a solution would be an adjusted set of prices $p$ that equal core payments $\bar{\rho}_{i}$ for each $i$ while maintaining WE, and are made by a valid cut $(\alpha,\alpha x^{*})$, according to the three sets of linear constraints, respectively. When this formulation is infeasible, its dual must be unbounded, noting that it is feasible when all variables are zero. This dual can be written:
Together with (13) and (14), we have:
$-\sum_{k\in\mathcal{K}^{*}}\bar{\rho}_{i(k)}y_{k}\geq\sum_{k\in\mathcal{K}\setminus\mathcal{K}^{*}}(b^{k}-b^{i(k)*}+\bar{\rho}_{i(k)})z_{k}$
@@
VCG payments can be linearized to PME prices if and only if VCG $=$ MRC.
@@
This induces $p=[0,20,15]$ for left pair or $p=[0,35,20]$ for the right pair, both resulting in linearizations of $\rho^{VCG}=[55,55]$. The left pair can be seen to price pairs of $A$ plus “only two of bids 1, 2, 3 can win,” while the right pair says more symmetrically, “only two of bids 1, 2, $k$ can win” for $k=3,4$. Notice that the top left cut is the sum of the right two. In our elaboration of MAP prices below, we will find minimal cuts that maintain balance among price-setting scenarios, and so we will favor the right two. Final cuts in our procedure belong to the Hilbert basis of the cone of feasible cuts, and so the top left cut could not be selected as an explanation of final payments.
@@
$\begin{array}[]{c}A\\ B\\ C\\ D\\ \end{array}$ $\left|\begin{array}[]{cc}\bm{8}&0\\ 8&0\\ 0&\bm{8}\\ 0&8\\ \end{array}\right|$ $\begin{array}[]{c}A\\ B\\ C\\ D\\ \end{array}\left|\begin{array}[]{cc}6&0\\ 0&\bm{6}\\ 2&0\\ 0&0\\ \end{array}\right|$ $\begin{array}[]{c}A\\ B\\ C\\ D\\ \end{array}\left|\begin{array}[]{cc}0&0\\ 2&0\\ 0&6\\ \bm{6}&0\\ \end{array}\right|$ (39)
@@
Figure 3: The core (in payment space) for Example 6.12. The pay-as-bid, NWE, and VCG $=$ MRC points are labeled. PME points form the shaded triangle on the near face, with the unique MAP point $\bm{(12,4,4)}$ in red.
@@
The MAP outcome for Example 6.12 can be found after adding two cuts, that say in words, respectively, “Bidders 1 and 2 cannot both get a most valued bundle,” and “Bidders 1 and 3 cannot both get a most valued bundle.” These cuts yield $p=[4,4,4,4,2,2]$ (the last two prices on artificial items) with total payments of 20 and bidder 1 paying for both artificial items. These payments $\rho_{2}=\rho_{3}=4$ lie strictly between the other outcomes for bidders 2 and 3, but with $\rho_{1}=12$ in all three outcomes. The core of Example  6.12 is shown in Figure 3, including the unique MAP point in red.
6.3 Computing MAP prices
We now introduce “Minimally Artificial PME” (MAP) payments as a precise solution concept. Our computational strategy is to find a minimal vector $\bar{\alpha}$ of artificial adjustments for each bid (positive real numbers), which determines final payments. Then in a second procedure, we produce an elaboration of these amounts into more meaningful prices, generated as a positive linear combination of minimal cut vectors. The first optimization discovers how much artificiality must be introduced, and the second elaboration operation finds a simplest set of cuts that explains the artificial amounts.
The first formulation (Find-$\bar{\alpha}$) minimizes the total amount of artificial (i.e., not linear in original auction items) adjustments needed to achieve PME:
$\min\quad$ $\bar{\alpha}\bm{1}$ (Find-$\bar{\alpha}$)
s.t. $\bar{\alpha}\bm{D}\geq\bm{0}$ (40)
$p\bm{A}+s\bm{B}+\bar{\alpha}\geq b$ (41)
@@ 6.3 Computing MAP prices
$\bar{\alpha}x^{-i}=\bar{\alpha}x^{*}$ $\forall i\in\mathcal{I}^{*}$ (44)
$pa^{k}+\bar{\alpha}_{k}\leq b_{k}\text{\quad\qquad when }x_{k}^{-i}=1$ $\forall i,k\in\mathcal{I}^{*}\times\mathcal{K}$ (45)
$\bar{\alpha}\in\mathbb{R}^{K}_{\geq 0}$ (46)
@@
Here we present Find-$\bar{\alpha}$ compactly, with nonlinear dependencies and quadratic constraints. An equivalent (but longer) linear integer program is presented in Appendix 9 for completeness.
@@
If there exists a NWE that is a PME, standard linear item prices are sufficient, and the objective will be zero, with $\bar{\alpha}=\bm{0}$, and thus MAP $=$ NWE. If, on the other extreme, all original items have excess supply, linear prices cannot be used at all, and $\bar{\alpha}$ does all the work of pricing with its nonlinear terms, as occurs in Examples 6.2 and 6.11.
@@
Constraints (40) like (10) ensure a valid cut, while (41) like (5) are WE-defining envy-freeness constraints, but here with the assistance of additional nonlinear $\bar{\alpha}$ terms. Constraints (42) and (43) ensure that when any winning bidder $i$ is removed, $x^{-i}$ variables select a feasible WDP solution without $i$, one that sells any item that receives a positive price. Set (44) ensures that (continuous) artificial item $\alpha$ sells completely in each $x^{-i}$ scenario. Together with (45) that ensure IR in each case, (42)-(45) explicitly fulfill the definition of PME when each bidder is removed. Surplus variables $s$ and thus (47) can be substituted out because $x^{*}$ is known, but we leave them in this formulation to show the connection to the simple WDP-dual. The complementary slackness condition shown in (48) is easy to implement simply as $p_{j}=0$ if item $j$ has excess supply under $x^{*}$.
@@
According to Theorem 6.3 and Corollary 6.8, formulation Find-$\bar{\alpha}$ will always have a solution, given the assumptions of this paper (a single-seller CA normalized to zero reserve prices). If these continuous MAP payment amounts $\bar{\alpha}$ are not all zero, we next perform what we will call elaboration to produce MAP item prices. (If $\bar{\alpha}=\bm{0}$ no elaboration is needed.) To do so, we first introduce some additional nomenclature.
[…]
$\bar{\bm{D}}_{k,\ell}=\begin{cases}\quad\text{ }1&\text{if }x^{\ell}_{k}=0\text{ , }x^{*}_{k}=1\text{ , and }\bar{\alpha}_{k}>0\\ \quad-1&\text{if }x^{\ell}_{k}=1\text{ , }x^{*}_{k}=0\text{ , and }\bar{\alpha}_{k}>0\\ \quad\text{ }0&\text{otherwise}\end{cases}$
That is, winning bids not in scenario $\ell$ get a positive 1 while non-winning bids in $\ell$ get -1, and the corresponding $\bar{\alpha}$ amounts of these two groups exactly balance for each $\ell$ column. (Winning bids that are in scenario $\ell$ effectively cancel out to a zero entry.) Note that the scenarios captured in the columns of $\bar{\bm{D}}$ are payment-setting to achieve PME by definition. For example, in Example 6.2, where “any two winners’ payments match the combined bids 4 and 5 which total 115,” the columns in $\bar{\bm{D}}$ correspond to the three (symmetric) scenarios in which one winner could join bids 4 and 5 as matching the total payments of all winners.
Under $\bar{\alpha}\neq\bm{0}$, at least one equation from (44) is non-zero, and so $\bar{\bm{D}}$ is not empty. From $\bar{\alpha}\bar{\bm{D}}=\bm{0}$ with nonzero $\bar{\alpha}$, $ker(\bar{\bm{D}}^{T})=\{\alpha|\alpha\bar{\bm{D}}=\bm{0}\}$ is nontrivial, i.e., $nullity(\bar{\bm{D}}^{T})=dim(ker(\bar{\bm{D}}^{T}))\geq 1$, from elementary linear algebra. This $nullity$ will be the dimension of our cut space (i.e., the number of cuts produced).
Definition 6.14
Critical bids are those appearing in at least one exact payment-matching scenario, and so must (critically) not see a total payment change during elaboration. Again, assuming an elaboration is needed, this set will be non-empty. The elaboration procedure will generate prices that maintain exactly minimal adjustments $\bar{\alpha}_{k}$ for all $\bar{\mathcal{K}}$ bids. Notice that $\{k|k\in\mathcal{K}^{*}\text{ and }\bar{\alpha}_{k}>0\}\subset\bar{\mathcal{K}}$ by (44).
Note that not all bids $k$ with $\bar{\alpha}_{k}>0$ are in $\bar{\mathcal{K}}$. That is, there are losing bids for which natural linear prices alone cannot satisfy envy-freeness, but which are not strong enough to help form a payment-matching scenario:
Definition 6.15
The set of non-critical bids is $\mathcal{K}^{nc}=\{k\in\mathcal{K}|\bar{\alpha}_{k}>0\text{ and }\forall\ell\text{ }\bar{\bm{D}}_{k,\ell}=0\}$.
It is not essential (and we find not desirable for interpretability) for non-critical bids to strictly adhere to their minimal adjustment amounts $\bar{\alpha}_{k}$, and so the elaboration procedure does not enforce this. (Notice also that there may be a third set of bids, $\mathcal{K}\setminus(\bar{\mathcal{K}}\cup\mathcal{K}^{nc})$, those bids with $\bar{\alpha}_{k}=0$. These bids satisfy envy-freeness via natural items alone, and may be referred to as irrelevant to the elaboration procedure.)
To form an elaboration of $\bar{\alpha}$ into a set of interpretable cuts, name this set of artificial cuts $\mathcal{J}^{+}$, with $J^{+}=|\mathcal{J}^{+}|$. These newly-added artificial items (cuts) $\alpha^{j}\in\mathcal{J}^{+}$ extend $\mathcal{J}$ (the natural items) and are generated to be concatenated to $\bm{A}$ as new rows, with prices $p_{j}$ for $j\in\mathcal{J}^{+}$ concatenated to $p$.
Definition 6.16
A set of vectors $\alpha^{j}$ for $j\in\mathcal{J}^{+}$ is an elaboration of $\bar{\alpha}$ if and only if $\exists p\in\mathbb{R}^{J^{+}}_{\geq 0}$ with:
$\sum_{j\in\mathcal{J}^{+}}p_{j}\alpha^{j}_{k}$ $=\bar{\alpha}_{k}$ $\forall k\in\bar{\mathcal{K}}$ (50)
$\sum_{j\in\mathcal{J}^{+}}p_{j}\alpha^{j}_{k}$ $\geq\bar{\alpha}_{k}$ $\forall k\in\bar{\mathcal{K}}^{nc}$ (51)
$\alpha^{j}\bm{D}$ $\geq\bm{0}$ $\forall j\in\mathcal{J}^{+}$ (52)
$\alpha^{j}$ $\in\mathbb{Z}^{K}_{\geq 0}$ $\forall j\in\mathcal{J}^{+}$ (53)
Here, (50) define a linear decomposition for critical bids, and (51) relax this decomposition for non-critical bids. The direction of this relaxation ensures envy-freeness is maintained; a non-critical bid will remain envy-free with prices totalling more than $\bar{\alpha}_{k}$. Critical bids, on the other hand, (those in $\bar{\mathcal{K}}$) match total payments of winning bids or are winning, and so strict equality is required. Constraints (52) assure that each element of the elaboration is a valid cut to WDP, and (53) assures that each artificial cut item consists of non-negative integer amounts in each bundle bid $k$, emulating natural-item supply constraints (2).
Proposition 6.17
Proof 6.18
$\min\quad$ $\alpha^{1}\bm{1}$ (Find-$\alpha^{1}$)
$\alpha^{1}\bar{\bm{D}}=\bm{0}$ (54)
$\alpha^{1}\bm{D}\geq\bm{0}$ (55)
$\alpha^{1}\bm{1}\geq 1$ (56)
$\alpha^{1}\in\mathbb{Z}^{K}_{\geq 0}$ (57)
In order to make progress with each basis cut added, we evaluate how much of $\bar{\alpha}$ can be covered by bases generated thus far by solving an LP, starting at $t=1$:
$\min\quad$ $r^{t}\bm{1}$ (PriceCutst)
$\sum_{j=1}^{t}p^{t}_{j}\alpha^{j}+r^{t}$ $=\bar{\alpha}$ (58)
$p^{t}$ $\in\mathbb{R}^{t}_{\geq 0}$ (59)
$r^{t}$ $\in\mathbb{R}^{K}_{\geq 0}$ (60)
Here $r$ is a vector of remainders, non-negative amounts of $\bar{\alpha}$ that cannot be covered by price-multiples of the $t$ cut vectors generated thus far. When at some iteration $t$ we arrive at zero objective (i.e., $r=\bm{0}$), part (i) of elaboration can terminate, having successfully found basis vectors of $ker(\bar{\bm{D}}^{T})$ that satisfy (50), (52), and (53). For example, if $nullity(\bar{\bm{D}}^{T})=1$, as the lone basis vector of $ker(\bar{\bm{D}}^{T})$, $\alpha^{1}$ would generate any kernel element as a multiple, and thus the optimal solution to PriceCutst is zero at the first iteration. If not, it is not time to terminate, and so we increment $t$ and generate the next cut vector by solving the following MIP (Find-$\alpha^{t}$) for $t\geq 2$:
$\min\quad$ $\alpha^{t}\bm{1}$ (Find-$\alpha^{t}$)
$\alpha^{t}+\sum_{j=1}^{t-1}\gamma_{j}\alpha^{j}=\delta^{t}$ (61)
$\alpha^{t}\bar{\bm{D}}=\bm{0}$ (62)
$\alpha^{t}\bm{D}\geq\bm{0}$ (63)
$\delta_{\tilde{k}}\geq\epsilon$ for some $\text{for some }\tilde{k}\in\bar{\mathcal{K}}\text{ with }r^{t-1}_{k}>0$ (65)
$\alpha^{t}\in\mathbb{Z}^{K}_{\geq 0}$ (67)
$\gamma\in\mathbb{R}^{t-1}$ (68)
$\delta^{t}\in\mathbb{R}^{K}_{\geq 0}$ (69)
Proposition 6.19
Proof 6.20
First, Find-$\alpha^{t}$ is feasible for $t\in\{1,...J^{+}\}=\mathcal{J}^{+}$. The feasible set $\{\alpha\in\mathbb{Z}^{K}_{\geq 0}|\alpha\bar{\bm{D}}=\bm{0},\alpha\bm{D}\geq\bm{0}\}$ consists of the integer lattice points of a pointed polyhedral cone, and thus is finitely generated by a unique Hilbert basis (see Aardal et al. 2002). This basis is a collection of component-wise minimal row vectors, together forming a matrix $H$ generating the same feasible set as $\{\alpha|\alpha=pH,p\in\mathbb{Z}^{|H|}_{\geq 0}\}$, which can also be characterized as a Graver basis (see Le and Römer 2024). The theoretical existence of this Hilbert basis guarantees integer solutions to each Find-$\alpha^{t}$. Because the Hilbert basis definition demands integer weights, $|H|>nullity(\bar{\bm{D}})$ is possible. Here, these weights are prices and thus continuous, and so a standard linear basis of size $nullity(\bar{\bm{D}})$ is appropriate. As a consequence, the generated $\alpha^{t}$ cuts are potentially a strict subset of $H$: generators of the standard linear cone generated by $H$. Note that there is no loss in restricting from the integer cone $\{\alpha\in\mathbb{Z}^{K}_{\geq 0}|\alpha\bar{\bm{D}}=\bm{0},\alpha\bm{D}\geq\bm{0}\}$ to the linear cone of $H$, as $\bar{\alpha}$ belongs to the latter, as it is rational given rational coefficients to the original problem, and therefore an integer multiple belongs to the former.
To see that progress in generating bases is made at each $t$, multiply (61) by an arbitrary $p_{t}>0$ and add it to the $t-1$ indexed (58), yielding:
$p_{t}\alpha^{t}+\sum_{j=1}^{t-1}(p_{t}\gamma_{j}+p^{t-1}_{j})\alpha^{j}+r^{t-1}-p_{t}\delta^{t}=\bar{\alpha}\text{ .}$
This new $\alpha^{t}$ extends the basis to improve the objective of PriceCutst, because $r^{t-1}-p_{t}\delta^{t}\leq r^{t-1}$ by construction, with strict inequality for the chosen $\tilde{k}$. This improvement would not be possible if $\alpha^{t}$ was a linear combination of previous $\alpha^{j}$s (or else optimality at $t-1$ is contradicted), thus the basis must grow. Note that this only shows that the new basis establishes an improving direction when $p_{t}$ is small enough that $p_{t}\gamma_{j}+p^{t-1}_{j}>0$ for $j<t$, and so we still solve PriceCutst in between each Find-$\alpha^{t}$ iteration.
$\min\quad\sum_{j\in\mathcal{J}^{+}}$ $\sum_{k\in\mathcal{K}^{nc}}\alpha^{j}_{k}$ ($nc$-lift)
$\sum_{j\in\mathcal{J}^{+}}$ $p_{j}\alpha^{j}_{k}\geq\bar{\alpha}_{k}$ $\forall k\in\mathcal{K}^{nc}$ (70)
To see that this this IP must be feasible, observe that we could selectively increase both bid amount $b^{k}$ and $\bar{\alpha}_{k}$ amounts by the same level for some bids $k\in\mathcal{K}^{nc}$ to result in a new auction in which all such $k$ are now critical. In that new auction, elaboration part (i) would give us cuts including $\mathcal{K}^{nc}$, and thus a feasible solution to $nc$-lift. Still, such an adjusted auction approach would tend to overstate the prices quoted to $nc$ bids and distort the interpretability of artificial items. We thus pursue the current formulation to find more informative prices.
  • Minimal: Implement elaboration exactly as stated here. For a selected final outcome based on a particular $\bar{\alpha}$. Non-critical bids are guaranteed to be among the AWE that minimize the total artificial price quotes to $nc$ bids.
  • $nc$-Nonlinear Only: Implement elaboration without running part (ii). Final prices are totally explained by the relationship to critical bids, including those that form the price-match allocations. Critical bids see the linear decomposition of $\bar{\alpha}_{k}=\sum_{j\in\mathcal{J}^{+}}p_{j}\alpha^{j}_{k}$, as always, while $nc$ bids would only see a lump-sum non-linear payment component $\bar{\alpha}_{k}$ (above linear prices on original items). This lump-sum would cover envy-freeness, and the optimality of Find-$\bar{\alpha}$ would still guarantee that the $nc$ bid could not be part of any allocation that would match total payments of the winners.

### v2

@@ 6.1 Cuts for WE
where $x^{f}$ is a fixed fractional solution to WDP-LP, treated here as a constant, as is $\bm{D}$. (If WDP-LP admits an integral solution, we are at a WE, and so we would terminate without the need for a new cut.) While constraints (10) ensure that $(\alpha,\alpha_{0})$ is a valid cut for WDP, any positive scalar multiple is also valid, resulting in an unbounded objective. To avoid this unboundedness, one can begin with $\alpha_{0}=1$ and later scale ($\alpha,\alpha_{0}$) to find an integer cut, as described in Algorithm 1, shown in Appendix 11. In practice, we do not need to construct the full $D$ matrix in order to solve this problem; we provide details on its dynamic generation in Appendix 10.
@@ 6.2 Cuts for any Core Point and PME
Example 4.2 demonstrated that a WE may not be a PME, motivating us to consider adding additional cuts to the WDP to get to PME. Algorithm 1 (see Appendix 11) can take us no further in generating new cuts once we have an AWE (or if we start at NWE), as its progress was driven by closing an IP gap that is now closed. For another, perhaps more interesting illustration, consider the single-minded bidders in the Example 6.1 that follows. Here there are $c_{A}=17$ copies of one item, $A$. (This problem is adapted from an example on lifting to get cuts to the knapsack problem from Bertsimas and Weismantel 2005.)
@@
VCG payments can be linearized to PME prices if and only if VCG == MRC.
@@
This induces $p=[0,20,15]$ for left pair or $p=[0,35,20]$ for the right pair, both resulting in linearizations of $\rho^{VCG}=[55,55]$. The left pair can be seen to price pairs of $A$ plus “only two of bids 1, 2, 3 can win,” while the right pair says more symmetrically, “only two of bids 1, 2, $k$ can win” for $k=3,4$. Notice that the top left cut is the sum of the right two. In our elaboration of MAP prices below, we will find minimal cuts that maintain balance among price-setting scenarios, and so we will favor the right two. Final cuts in our procedure belong to the Hilbert basis of the cone of feasible cuts, and so the top left cut could not be selected as an explanation of final payments. (The Hilbert basis is the unique set of minimal integer cuts from which any other integer cut could be generated through non-negative integer combination. See Aardal et al. (2002), for example.)
@@
[…]
@@
Figure 3: The core (in payment space) for Example 6.9. The pay-as-bid, NWE, and VCG == MRC points are labeled. PME points form the shaded triangle on the near face, with the unique MAP point $\bm{(12,4,4)}$ in red.
@@
The MAP outcome for Example 6.9 can be found after adding two cuts, that say in words, respectively, “Bidders 1 and 2 cannot both get a most valued bundle,” and “Bidders 1 and 3 cannot both get a most valued bundle.” These cuts yield $p=[4,4,4,4,2,2]$ (the last two prices on artificial items) with total payments of 20 and bidder 1 paying for both artificial items. These payments $\rho_{2}=\rho_{3}=4$ lie strictly between the other outcomes for bidders 2 and 3, but with $\rho_{1}=12$ in all three outcomes. The core of Example 6.9 is shown in Figure 3, including the unique MAP point in red.
7 Computing MAP prices
Our computational strategy is to find a minimal vector $\phi$ of artificial adjustments for each bid (positive real numbers), which determines final payments. Then in a second procedure, we produce an elaboration of these amounts into more meaningful prices, generated as a positive linear combination of minimal cut vectors. The first optimization discovers how much total artificiality must be introduced, and the second elaboration operation finds a simplest set of cuts that explains the artificial amounts.
The first formulation (min-$\phi$) minimizes the total amount of artificial (i.e., not linear in original auction items) adjustments needed to achieve PME:
$\min\quad$ $\phi\bm{1}$ (min-$\phi$)
s.t. $\phi\bm{D}\geq\bm{0}$ (37)
$p\bm{A}+s\bm{B}+\phi\geq b$ (38)
@@ 7 Computing MAP prices
$\phi x^{-i}=\phi x^{*}$ $\forall i\in\mathcal{I}^{*}$ (41)
$pa^{k}+\phi_{k}\leq b_{k}\text{\quad\qquad when }x_{k}^{-i}=1$ $\forall i,k\in\mathcal{I}^{*}\times\mathcal{K}$ (42)
$\phi\in\mathbb{R}^{K}_{\geq 0}$ (43)
@@
Here we present min-$\phi$ compactly, with nonlinear dependencies and quadratic constraints. An equivalent (but longer) linear integer program is presented in Appendix 12 for completeness.
@@
If there exists a NWE that is a PME, standard linear item prices are sufficient, and the objective will be zero, with $\phi=\bm{0}$, and thus MAP == NWE. If, on the other extreme, all original items have excess supply, linear prices cannot be used at all, and $\phi$ does all the work of pricing with its nonlinear terms, as occurs in Examples 6.1 and 6.8.
@@
Constraints (37) like (10) ensure a valid cut, while (38) like (5) are WE-defining envy-freeness constraints, but here with the assistance of additional nonlinear $\phi$ terms. Constraints (39) and (40) ensure that when any winning bidder $i$ is removed, $x^{-i}$ variables select a feasible WDP solution without $i$, one that sells any item that receives a positive price. Constraints (41) ensures that (continuous) artificial item $\phi$ sells completely in each $x^{-i}$ scenario. Together with (42) that ensure IR in each case, (39)-(42) explicitly fulfill the definition of PME when each bidder is removed. Surplus variables $s$ and thus (44) can be substituted out because $x^{*}$ is known, but we leave them in this formulation to show the connection to the simple WDP-dual. The complementary slackness condition shown in (45) is easy to implement simply as $p_{j}=0$ if item $j$ has excess supply under $x^{*}$.
@@
According to Theorem 6.2 and Corollary 6.5, formulation min-$\phi$ will always have a solution, given the assumptions of this paper (a single-seller CA normalized to zero reserve prices). If these continuous MAP payment amounts $\phi$ are not all zero, we next perform what we will call elaboration to produce MAP item prices. (If $\phi=\bm{0}$ no elaboration is needed.) The elaboration proceeds using a sequence of IPs, each generating a single artificial item (cut), resulting in a set of minimal integral cuts as shown in each of the illustrative examples throughout the paper. Details of this procedure are given in Appendix 13, which can be summarized as iteratively generating a minimal linear basis for the null space of the tight $\bm{D}$ constraints in formulation min-$\phi$.
For instance, consider again Example 6.1. Because only 15 of the 17 copies of natural item $A$ sell, $p_{A}=0$ (by definition of a WE), and all pricing must be attributed to artificial items. The solution to min-$\phi$ is $\phi=(57.5,57.5,57.5,25,90)$, with all tight $\bm{D}$ constraints saying that the $\phi$ amounts of any two winning bids (from $\{1,2,3\})$ must match the losing bids’ $\phi$ amounts (from $\{4,5\}$), for example, $\phi_{1}+\phi_{3}\geq\phi_{4}+\phi_{5}$. Elaboration looks first for a valid cut with minimal total integer coefficients, breaking ties arbitrarily. Suppose it first finds $x_{1}+x_{2}+x_{3}+x_{4}+x_{5}\leq 3$, or extracting just LHS coefficients, $\alpha^{1}=(1,1,1,1,1)$, a Hilbert basis of the cone of tight cuts. The procedure next attempts to generate $\phi$ as a positive multiple, but cannot do so without a positive remainder called $r^{t}$ at iteration $t$ in the Appendix 13 notation. So far we have $\phi=p^{1}_{1}\alpha^{1}+r^{1}=25(1,1,1,1,1)+(32.5,32.5,32.5,0,65)$. ‘1Iterating, a next Hilbert basis element gives cut $x_{1}+x_{2}+x_{3}+x_{5}\leq 3$ or $\alpha^{2}=(1,1,1,0,2)$. The elaboration algorithm can now terminate, as these two basis elements can decompose $\phi$ with zero remainder: $\phi=p^{2}_{1}\alpha^{1}+p^{2}_{2}\alpha^{2}+r^{2}=25(1,1,1,1,1)+32.5(1,1,1,0,2)+\bm{0}$.